# Meritocracy and Its Discontents: Long-run Effects of Repeated School Admission Reforms

### v1

@@ 2.1 Bidirectional Admission Reforms
This institutional innovation was short-lived, however. Due to political and administrative reasons detailed below, the government switched back to Dapp (with a unified exam) in 1908. The government then continued to oscillate between decentralization and centralization, reintroducing Capp in 1917, moving back to Dapp (with a unified exam) in 1919, reinstituting Capp (with major modifications of allowing applicants to list at most two schools) in 1926, and finally settling down to Dapp (with separate exams) in 1928. In a space of thirty years, there were three periods of centralized admissions: first in 1902–1907, next in 1917–1918, and finally in 1926–1927.191919For details of the evolution of the admission system, see Appendix Table A.1.
We exploit the series of bidirectional policy changes to examine the impacts of centralization. Importantly, it was difficult for applicants to anticipate the exact timing of these policy changes, as the admission policy was decided and announced only a few months in advance. This suggests that the reform timing is close to exogenous, as we further explore in Section 3.5.
@@ 2.2 Predicted Impacts
To guide our empirical analysis, we first predict effects of centralization by using historical documents and theoretical insights. Historical documents suggest that the repeated policy changes resulted from intense bargaining between the Ministry of Education in the central government and the Council of School Principals representing local interests. [^20] Detailed accounts are provided in Moriguchi (2021), pp.195-199. While the Ministry of Education emphasized the importance of using centralized admissions to enroll the best and brightest to the national higher education, the Council of School Principals was opposed to centralization, pointing out several adverse consequences for rural schools and their local communities.
@@
First, according to the Council of School Principals, the centralized system encourages applicants to be more risk-taking and rank the most competitive and prestigious urban school as the first choice. This is because the centralized system provides applicants with options to apply for lower choice schools as insurance.
Hypothesis 1. Under the centralized system, more applicants in all areas rank most prestigious School 1 as their first choice.
@@
Second, Hypothesis 1 means that the centralized system made a greater number of rural applicants apply to and enter urban schools. In addition, the centralization allowed high-achieving applicants in urban areas to enter rural schools even if they failed at the most selective urban school. This, in turn, often resulted in the rejection of lower-achieving rural applicants who would have been able to enter their local schools under decentralized admissions. These effects would jointly increase the mobility of students across the country.
Hypothesis 2. Under the centralized system, a smaller share of each school’s seats are allocated to its local students.
The Council of School Principals also complained that the centralized system lets many reluctant and unmotivated students enter rural schools as fallback options, which lowered school morale.
Third, by the meritocracy constraint embedded in the centralized system, it allocated a higher fraction of seats at Schools 1–8 to high-achieving applicants. After reviewing the admission results of 1917 (the first year of the second centralization period), the Council of School Principals found that the number of rural middle-school graduates not admitted to any of Schools 1–8 increased considerably compared to the previous three years of decentralized admissions. This is presumably because high-achieving applicants were located disproportionately in urban areas.
Hypothesis 3. Under the centralized system, a larger share of all school seats in Schools 1–8 are allocated to urban-born applicants.
Online Appendix B formalizes these hypotheses as theoretical predictions in particular models. A noted historian summarizes these observations as follows (Takeuchi, 2011, p.121):
Urban applicants ‘overwhelm’ rural applicants by applying for rural schools as fallback options. Urban applicants rob rural applicants of opportunities that were once open to them. This ruins the meaning of building national higher schools across the nation.
This upset local communities of rural schools, especially because they contributed land and other resources to the construction of rural schools (Takeuchi, 2011, p.56 and pp.106-107). [^21] An additional concern against the centralized system was the administrative cost of implementing it in the absence of modern computers and photocopying technologies. Both the Ministry of Education officials and the school principals repeatedly pointed out the difficulty of grading thousands of exam sheets by hand in a short period of time and assigning applicants to schools according to the algorithm (Takeuchi, 2011, pp.118-119; Moriguchi, 2021, p.199). The school principals also deemed centralization as an intrusion on their power and autonomy.
These documents highlight a potential meritocracy-equity tradeoff of centralization. On the one hand, the centralized admissions made the school seat allocation more meritocratic, enabling high-achieving students to enter one of the elite schools even if they failed at the most selective one. On the other hand, this meritocracy might have come at the expense of equal regional access to higher education, whereby high-achieving urban applicants dominated rural applicants. Below we test the above hypotheses as well as their persistence in the long run.
3 Short-run Impacts
@@ 3.1 Data
To analyze short-run effects of the centralized admission system, we assemble and digitalize data on application and enrollment. First, we collect data on the number of applicants by school of their first choice from 1900 to 1930, using Ministry of Education Yearbook and other sources. For these two years (1916 and 1917), we also collect data on the number of applicants by school of their first-choice and by prefecture of their middle school. Appendix Section A.3.1 provides detailed explanations for data sources and definitions.
@@ 3.2 Strategic Responses by Applicants
As an immediate effect, switching back and forth between the centralized and the decentralized admission systems causes stark strategic responses in application behavior. Figure 1(a) shows that the three periods of Capp are associated with sharp increases in the share of applicants who chose the most selective School 1 in Tokyo as their first choice, as predicted.
To observe regional variations, we examine how the propensity of applicants to rank School 1 as their first choice changes between 1916 (under Dapp) and 1917 (under Capp). The results in Appendix Table A.3 show that the meritocratic centralization induced applicants in all regions to rank the most prestigious school first and to make more long-distance applications. As a result, the competition to enter School 1 became even more intense under Capp (Appendix Figure A.4). As a result, under Capp, only a small fraction of first-choice applicants were admitted to School 1, producing many high-scoring applicants who were rejected by School 1 and then assigned to their second or lower-choice school.
@@ 3.3 Regional Mobility in Enrollment
The changes in the mechanism and behavior influence enrollment outcomes. To measure the geographical mobility of entrants, we compute the “enrollment distance” defined as the direct distance between an entrant’s birth prefecture and the school he entered. As Figure 1(b) shows, the centralized system is associated with a sharp and discontinuous increase in enrollment distance. [^22] The effect is especially stark in the first two periods of Capp. The third period of Capp in 1926–27 was qualitatively different from that in the first and second periods. In the third period of Capp, schools were divided into two groups and applicants were allowed to choose and rank at most two schools (one school per group) in 1926–27. This increase in regional mobility is driven by a sharp reduction in the number of “local entrants” defined as entrants who entered a school in (or near) their birth prefecture. To show this, we estimate the following regression for each school $s$ separately:
$\begin{split}Y_{pt}^{s}=&\ \beta_{1}^{s}\times Centralized_{t}\times 1\{\text{school }s\text{ is located in prefecture }p\}\\ &+\beta_{2}^{s}\times Centralized_{t}\times 1\{\text{school }s\text{ is 1-100 km away from prefecture }p\}\\ &+\alpha^{s}X_{pt}+\gamma_{t}^{s}+\gamma_{p}^{s}+\epsilon_{pt}^{s},\end{split}$ (1)
@@
As shown in Panel (a) of Table 1, Capp reduces the number of local entrants born in the school’s prefecture across the country. The coefficients $\beta_{1}^{s}$ are significantly negative for all schools. Column (1) shows that the number of School 1 entrants born in Tokyo prefecture declined by about 25%. Schools 4–7 experienced reductions in the number of entrants born not only from the school’s prefecture but also from surrounding prefectures. In other words, centralization weakened the local monopoly power of each school by creating a national market for higher education, consistent with our theoretical prediction in Appendix B (Proposition 3).
3.4 Meritocracy vs Equal Regional Access
@@
Which regions gained more school seats under the centralized system? We now investigate how centralization altered the geographical allocation of school seats. Figure 2(a) plots the change in the total number of entrants to Schools 1–8 from Dapp to Capp by their birth prefecture, where the darker blue color indicates the greater decline and the darker red color indicates the greater increase. The figure shows that Tokyo prefecture (where School 1 was located) and its surrounding area gained school seats under Capp, while most of western and northern prefectures lost school seats.
@@
To see this quantitatively, we regress the number of entrants to Schools 1–8 in year $t$ born in prefecture $p$ on the interaction terms between $Centralized_{t}$ and each of the following two indicator variables: the indicator that takes 1 if prefecture $p$ was Tokyo and the indicator that takes 1 if prefecture $p$ was 1-100 km away from Tokyo. To control for the decline of local entrants, we also add the interaction terms between $Centralized_{t}$ and each of the following two indicator variables: the indicator that takes 1 if any of Schools 1–8 was located in prefecture $p$ and the indicator that takes 1 if any of Schools 1–8 was located 1-100 km away from prefecture $p$. As in Equation (1), we also control for observable characteristics of prefecture $p$ and year $t$ and include year and prefecture fixed effects.
@@
Column (1) in Table 1 Panel (b) shows the number of entrants born in Tokyo prefecture increased by 13% per year under Capp relative to its mean under Dapp. The increase in the number of entrants is even higher for its surrounding area; the number of entrants born in prefectures 1-100 km away from Tokyo increased by 45% per prefecture per year under Capp relative to its mean under Dapp. [^23] One possible reason for this concentration of the effect in surrounding prefectures is that applicants from this area might be marginal applicants who would not be admitted under Dapp but would be admitted under Capp. On the other hand, many high-scoring applicants from the Tokyo prefecture might be admitted even under Dapp. Consistent with this hypothesis, the number of entrants from Tokyo (201) is much larger than the number of entrants from surrounding prefectures (26 per prefecture) under Dapp.
@@
Motivated by these observations, we define “Tokyo area” as the set of prefectures located within 100 km from Tokyo (see Appendix Table A.3 for its location). Figure 2(b) depicts the time evolution of the share of entrants to Schools 1–8 who were born in the Tokyo area. Consistent with the above results, the share of Tokyo-area born entrants rose during the periods of Capp.
@@
To see which schools received more Tokyo-area born entrants under Capp, we regress the number of entrants to each school $s$ in year $t$ born in prefecture $p$ on the interaction terms between $Centralized_{t}$ and the variables indicating whether entrants’ birth prefecture $p$ is Tokyo or near Tokyo. We include the same variables used in Equation (1) to control for school locations and other prefecture and year characteristics. Under Capp, a greater number of Tokyo-area-born students entered less selective rural schools, as shown in Columns (2)–(9) in Table 1 Panel (b). [^24] The results remain almost the same when we additionally control for observable prefecture characteristics (see Appendix Table A.4).
In summary, the net effect of Capp is such that the increased inter-regional applications caused high-achieving students in the Tokyo area to crowd out lower-achieving, rural-born students from their local schools. This evidence is consistent with the historian’s observation (Takeuchi (2011) cited in Section 2.2) that, as a result of the centralization, urban applicants mainly from Tokyo dominated rural applicants in entering national elite schools.
@@
The Tokyo area shares common characteristics with what is generally considered urban. Prefectures in the Tokyo area had a larger population, higher income per capita, and better educational infrastructure (indicated by a greater number of middle-school graduates eligible for higher education) compared to other prefectures (Appendix Table A.5). In Appendix Table A.6, we use these urban characteristics instead of the Tokyo area indicator in the same specification as Column (1) in Panel (b) of Table 1. The result shows that these urban characteristics, especially educational infrastructure, positively predict the areas which produced a greater number of entrants under Capp. Therefore, a greater number of high achieving students in the Tokyo area may explain why the Tokyo area gained more school seats under centralization.
@@
Students’ heterogenous preferences across areas can be another reason why the Tokyo area gained more school seats under the centralized system. Students prefer their local school for geographical and cultural proximity. The local school for students born in the Tokyo area is the most competitive School 1. Holding academic achievement constant, during the decentralization, students born in the Tokyo area are more likely to apply to School 1 and therefore fail to enter any school (even though they might have been able to enter Schools 2–8 had they applied). This hypothesis is consistent with Appendix Table A.6, which shows that the share of applicants to School 1 during the decentralization is associated with a greater number of entrants to Schools 1–8 during the centralization.
In summary, our analysis suggests that a combination of better educational infrastructure and strong preference for School 1 explains why the Tokyo area gained more school seats under the centralized system.
@@ 3.5 Other Institutional Changes
We discuss potential threats to our empirical analysis. In particular, if changes in other institutional and policy factors were correlated with the admission reforms, it could influence application behavior and enrollment outcomes, explaining our finding that more applicants applied for School 1 during centralization. Such other institutional factors include simultaneous reforms in middle schools, the total number of applicants, capacities at Schools 1–8, and the capacity of School 1 relative to the capacities of other schools.
We investigate these concerns and confirm that centralization periods are not correlated with time-series changes in the following variables (Columns (1)–(6) in Appendix Table A.7): the number of middle school graduates, the national number of applicants as well as the level of competitiveness (measured by the number of entrants divided by the number of applicants), the total number of entrants to Schools 1–8, the share of entrants to School 1 in all entrants, the probability of unsuccessful applicants retaking the exam in subsequent years, and the average age of entrants.
@@
Finally, the centralization reform introduced not only the meritocratic assignment algorithm, but also the unified entrance exam that applicants could take at any school location (see Appendix Table A.1). As such, the estimated impacts of centralization may be confounded by the unification of entrance exams and more flexible exam location choices. To investigate this issue, we analyze how key outcomes change from 1900 to 1901, during which the government also introduced a single entrance exam that applicants were allowed to take anywhere while the assignment method remained unchanged (i.e., decentralized). Figures 1(a) and 1(b) show that this institutional change from 1900 to 1901 induced little changes in application and enrollment patterns. The estimated impacts of centralization are therefore likely due to the meritocratic assignment algorithm rather than the changes in exam contents and locations.

### v2

@@ 2.2 Political Economy of the Admission Reforms
This institutional innovation was short-lived, however. Due to the opposition detailed below, the government switched back to decentralization (with a unified exam) in 1908. The government then continued to oscillate between the two regimes, reintroducing centralization in 1917, moving back to decentralization (with a unified exam) in 1919, reinstituting centralization (with modifications of allowing applicants to list at most two schools) in 1926, and finally settling down to decentralization (with separate exams) in 1928. In a space of thirty years, there were three periods of centralized admissions: first in 1902–1907, next in 1917–1918, and finally in 1926–1927.191919For details of the evolution of the admission system, see Appendix Table A.1. We exploit the series of bidirectional policy changes to examine the impacts of centralization.
Historical documents suggest that the repeated policy changes were a result of annual bargaining between the Ministry of Education (representing the central government) and the Council of School Principals (representing local interests of Schools 1–8). [^20] Detailed accounts are provided in Moriguchi (2021), pp.195-199. The admission policy, as a result of bargaining, was decided only a few months before the exam, making it difficult for applicants to anticipate the exact timing of the reforms. Indeed, we confirm that the timing of the reforms is not associated with the size and composition of applicants, school capacities, and other potential confounding factors (see Section 3.6).
Throughout the bargaining process, the Ministry of Education insisted on centralized admissions to select the best and brightest, while the Council of School Principals was opposed to centralization, pointing out its adverse impacts on rural schools located in remote regions. [^21] Other reasons against centralization included a loss of school autonomy and independence as well as the high administrative cost of implementing centralized admissions. They claimed that these rural schools lost the most talented students in their local areas (who would have entered local schools under decentralization) to urban schools, especially School 1. They also complained that rural schools instead received many reluctant and unmotivated students who only came to these schools as fallback options.
Moreover, a review of admission results revealed that the number of applicants from rural areas (areas where rural schools were located) not admitted to any of Schools 1–8 increased considerably under centralization. [^22] Ministry of Education (1917), Report on National Higher School Entrance Examination of 1917: Extra Issue. This was upsetting to rural schools and their local communities, as rural communities contributed land and other resources to construct rural schools. A noted historian summarizes the situation as follows (Takeuchi, 2011, p.121):
Urban applicants ‘overwhelmed’ rural applicants by applying for rural schools as fallback options. Urban applicants robbed rural applicants of opportunities that were once open to them. This ruined the meaning of building national higher schools across the nation.
These observations highlight a potential meritocracy-equity tradeoff of centralization. On one hand, the centralized admissions made the school seat allocation more meritocratic, enabling high-achieving students to enter one of Schools 1–8 even if they failed at the most prestigious one. On the other hand, this meritocracy might have come at the expense of equal regional access to national higher education, whereby high-achieving urban applicants dominated rural applicants.
3 Short-run Impacts
@@ 3.1 Predicted Impacts
Motivated by the above historical observations, we first make predictions on the short-run effects of centralization. Online Appendix B formalizes the predictions using a simple model (Propositions 1, 2, and 3).
@@
First, we predict changes in application behavior. Specifically, the centralized system would encourage applicants all over the country to rank the most preferred and selective urban school as their first choice. This is because the centralized system provides applicants with an option to apply for second and lower choice schools as insurance.
Second, if high-achieving applicants are disproportionately located in a specific geographical area, such as an urban area, then a larger fraction of all school seats in Schools 1–8 would be assigned to students coming from that area under the centralized system. This is because of the meritocratic constraint (Step 1) in the centralized assignment algorithm, which assigns only the top-scoring applicants to any school.
@@
Finally, a smaller fraction of applicants would be assigned to their local school under centralized admissions. Intuitively, meritocratic centralization enables high-scoring applicants from urban areas to enter rural schools even if they fail to enter the most competitive urban school. At the same time, the centralized system would make a greater number of high-scoring rural applicants apply to and enter urban schools. Under the centralized system, therefore, rural students would more often move to urban areas, while urban students would more often move to rural areas. These effects would jointly increase the spatial mobility of students.
@@ 3.2 Data
To empirically analyze short-run effects of the centralized admission system, we newly assemble data on application and enrollment. First, we collect data on the number of applicants by school of their first choice from 1900 to 1930, using the Ministry of Education Yearbook and other sources. For two specific years (1916 and 1917), we have additional data on the number of applicants by school of their first-choice and by prefecture of their middle school. Japan was divided into 47 prefectures. Appendix Section A.3.1 provides detailed explanations for data sources and definitions.
@@ 3.3 Strategic Responses by Applicants
As an immediate effect, switching back and forth between the centralized and the decentralized admission systems causes stark strategic responses in application behavior. Figure 1(a) shows that the three periods of centralization are associated with sharp increases in the share of applicants who chose the most selective School 1 in Tokyo as their first choice, as predicted.
To observe regional variations, we examine how the propensity of applicants to rank School 1 as their first choice changes between 1916 (under decentralization) and 1917 (under centralization). The results in Appendix Table A.3 show that meritocratic centralization induced applicants in all regions to rank the most prestigious school first and to make more long-distance applications. The competition to enter School 1 thus became even more intense under centralization (Appendix Figure A.4). As a result, under centralization, only a small fraction of first-choice applicants were admitted to School 1, producing many high-scoring applicants who were rejected by School 1 and then assigned to their second or lower-choice school.
@@ 3.4 Regional Mobility in Enrollment
The changes in the assignment algorithm and application behavior influence enrollment outcomes. To measure the geographical mobility of entrants, we compute the “enrollment distance” defined as the direct distance between an entrant’s birth prefecture and the school he entered. As Figure 1(b) shows, the centralized system is associated with a sharp and discontinuous increase in enrollment distance. [^23] The effect is especially stark in the first two periods of centralization. The third period of centralization in 1926–27 was qualitatively different from that in the first and second periods. In the third period of centralization, schools were divided into two groups and applicants were allowed to choose and rank at most two schools (one school per group) in 1926–27. This increase in regional mobility is driven by a sharp reduction in the number of “local entrants” defined as entrants who entered a school in (or near) their birth prefecture. To show this, we estimate the following regression for each school $s$ separately:
$\begin{split}Y_{pt}^{s}=&\ \beta_{1}^{s}\times Centralized_{t}\times 1\{\text{school }s\text{ is located in prefecture }p\}\\ &+\beta_{2}^{s}\times Centralized_{t}\times 1\{\text{school }s\text{ is 1--100km away from prefecture }p\}\\ &+\alpha^{s}X_{pt}+\gamma_{t}^{s}+\gamma_{p}^{s}+\epsilon_{pt}^{s},\end{split}$ (1)
@@
As shown in Panel (a) of Table 1, centralization reduces the number of local entrants born in the school’s prefecture across the country. The coefficients $\beta_{1}^{s}$ are significantly negative for all schools. Column (1) shows that the number of School 1 entrants born in Tokyo prefecture declined by about 26% relative to its mean under decentralization. Schools 2–7 experienced reductions in the number of entrants born not only from the school’s prefecture but also from surrounding prefectures. In other words, centralization weakened the local monopoly power of each school by integrating local markets into a national market, consistent with our prediction.
3.5 Meritocracy vs Equal Access
@@
We now investigate the distributional effect of the centralized system on the allocation of school seats at Schools 1–8. Figure 2(a) plots the change in the number of entrants to Schools 1–8 from decentralization to centralization by their birth prefecture, where the darker blue color indicates the greater decline and the darker red color indicates the greater increase. The figure shows that Tokyo prefecture (where School 1 was located) and its surrounding area gained school seats under centralization, while most of western and northern prefectures lost school seats.
@@
To see this quantitatively, we regress the number of entrants to Schools 1–8 in year $t$ born in prefecture $p$ on the interaction terms between $Centralized_{t}$ and each of the following two indicator variables: the indicator that takes 1 if prefecture $p$ was Tokyo and the indicator that takes 1 if prefecture $p$ was 1–100 km away from Tokyo. To control for the decline of local entrants, we also add the interaction terms between $Centralized_{t}$ and each of the following two indicator variables: the indicator that takes 1 if any of Schools 1–8 was located in prefecture $p$ and the indicator that takes 1 if any of Schools 1–8 was located 1–100 km away from prefecture $p$. As in Equation (1), we also control for observable characteristics of prefecture $p$ and year $t$ and include year and prefecture fixed effects.
@@
Column (1) in Table 1 Panel (b) shows the number of entrants born in Tokyo prefecture increased by 12% per year under centralization relative to its mean under decentralization. The increase in the number of entrants is even higher for its surrounding area; the number of entrants born in prefectures 1–100 km away from Tokyo increased by 45% per year under centralization relative to its mean under decentralization. [^24] One possible reason for this concentration of the effect in surrounding prefectures is that applicants from this area might be marginal applicants who would not be admitted under decentralization but would be admitted under centralization. By contrast, many high-scoring applicants from the Tokyo prefecture might be admitted even under decentralization. Consistent with this hypothesis, the number of entrants from Tokyo (201) is much larger than the number of entrants from surrounding prefectures (26 per prefecture) under decentralization.
@@
Motivated by these observations, we define the “Tokyo area” as the set of prefectures located within 100 km from Tokyo (see Appendix Table A.3 for its location). Figure 2(b) depicts the time evolution of the share of entrants to Schools 1–8 who were born in the Tokyo area. Consistent with the above results, the share of Tokyo-area-born entrants rose during the periods of centralization.
@@
To see which schools received more Tokyo-area-born entrants under centralization, we regress the number of entrants to each school $s$ in year $t$ born in prefecture $p$ on the interaction terms between $Centralized_{t}$ and the variables indicating whether entrants’ birth prefecture $p$ is Tokyo or near Tokyo. We include the same variables used in Equation (1) to control for school locations and other prefecture and year characteristics. Under centralization, more Tokyo-area-born students entered less selective rural schools, as shown in Columns (2)–(8) in Table 1 Panel (b). [^25] The results remain almost the same when we additionally control for observable prefecture characteristics (see Appendix Table A.4). In summary, the net effect of centralization is such that the increased inter-regional applications caused high-achieving, urban-born students to crowd out lower-achieving, rural-born students from their local schools.
@@
The Tokyo area shares common characteristics with what is generally considered urban and with high socioeconomic status. Prefectures in the Tokyo area had a larger population, higher income per capita, and better educational infrastructure (indicated by a greater number of middle-school graduates eligible for higher education) compared to other prefectures (Appendix Table A.5). In Appendix Table A.6, we use these urban characteristics instead of the Tokyo area indicator in the same specification as Column (1) in Panel (b) of Table 1. The result shows that these urban characteristics, especially educational infrastructure, positively predict the areas which produced a greater number of entrants under centralization. Therefore, richer educational infrastructures and a greater number of high achieving students in the Tokyo area may explain why the Tokyo area gained more school seats under centralization.
@@
Students’ heterogeneous preferences across school locations can be another reason why the Tokyo area gained more school seats under the centralized system. Students prefer their local school for geographical and cultural proximity. The local school for students born in the Tokyo area is the most competitive School 1. Holding academic achievement constant, during the decentralization, students born in the Tokyo area are more likely to apply to School 1 and therefore fail to enter any school (even though they might have been able to enter Schools 2–8 had they applied). This hypothesis is consistent with Appendix Table A.6, which shows that the share of applicants to School 1 during decentralization is associated with a larger increase in the number of entrants to Schools 1–8 during centralization.
In summary, the net effect of centralization is such that the increased regional mobility caused urban-born high achievers with high socioeconomic statuses to crowd out rural students from elite higher education. Key mechanisms behind this distributional effect are better educational infrastructures and strong preference for School 1.
@@ 3.6 Other Institutional Changes
We finish the short-run analysis by discussing potential threats. In particular, if changes in other institutional and policy factors were correlated with the admission reforms, it could influence application behavior and enrollment outcomes, explaining our finding that more applicants applied for School 1 during centralization. Such other institutional factors include simultaneous reforms in middle schools, the total number of applicants, capacities at Schools 1–8, and the capacity of School 1 relative to the capacities of other schools.
We investigate these concerns and confirm that centralization periods are not correlated with time-series changes in the following variables (Columns (1)–(6) in Appendix Table A.7): the number of middle school graduates, the national number of applicants as well as the level of competitiveness (measured by the number of entrants divided by the number of applicants), the total number of entrants to Schools 1–8, the share of entrants to School 1 in all entrants, the probability of unsuccessful applicants retaking the exam in subsequent years, the average age of entrants, and government expenditure on higher education.
@@
Finally, the centralization reform introduced not only the meritocratic assignment algorithm, but also the unified entrance exam that applicants could take at any school location (see Appendix Table A.1). As such, the estimated impacts of centralization may be confounded by the unification of entrance exams and more flexible exam location choices. To investigate this issue, we analyze how key outcomes change from 1900 to 1901, during which the government also introduced a single entrance exam that applicants were allowed to take at any school while the assignment method remained decentralized. Figures 1(a) and 1(b) show that this institutional change from 1900 to 1901 induced little changes in application and enrollment patterns. The estimated impacts of centralization are therefore likely due to the meritocratic assignment algorithm rather than the changes in exam contents and locations.